Back to the NIKA2 main page

Daily Reports

2024 Mar. 5th, Tuesday

Sofia, Frédéric, Tuan-Anh and Angel at the telescope since 10:30 (hereafter, local time).

16:00-16:50; Angel: Starting system check. It was necessary to restart the BluePill for moving the telescope. DAQ started at 16:30 (lost KIDs: 18, 1, 38). Pointing on Uranus (scan 113); imbFITS files properly created. Despite we tried to keep the main dish as far as possible from solar radiation during the maintenance slot, a focus sequence check gave a possible correction far from the expected: sign of a distorted beam. End of system check and observations suspended at 16:50: Aleksej K. (OHB) requires the telescope for some tests.

18:00-00:00; Sofia and Angel: Resuming observations after OHB's tests. After doing pointing and focus on source 0736+017, we had a beam of 17.5 x 18.5. Completed repetitions of WB89-670 (1 repetition done, 1 remaining). Then switched to project 151-23 targeting cluster PSZ2G210 (5th repetition out of 6, 40 min long). Waiting for a reasonable relaxation of the telescope. Very stable atmosphere conditions, peak-to-peak variations below 2 Jy/beam. Moved to 0316+413 for a beam map in total power (scan 157), after pointing and focus. Moved to calibrator CRL618, set pointing 7.1 8.1. FWHM 17.16 x 18.17, calib_1scan (scan 159) with nikaw-23. Performed a skydip (scan 160). Then moved to the cluster PSZ2G206 for project 151-23, new pointing on nearby source 0735+178, set pointing 5.4 7.6. Completing final repetition of PSZ2G206 (6/6).

2024 Mar. 6th, Wednesday

00:00-06:00; Frédéric: I started with a pointing and a focus on 1045+018 (FWHM 16.8"x18.3") under nikaw-23. At 01:37 (local time), I went to project 138-23 to observe NGC4214 (2 repetitions after pointing at 1156+295), before the source went above 80 degrees. I moved to project 150-23, while NGC4214 was at zenith. I observed id321 (1 repetition after pointing at 1044+719; FWHM 16.77"x18.14"). I launched another repetition on id321 (after pointing and focus on 1044+719; FWHM 16.7"x18.48") before it went under 40 degrees. At 05:27 (local time), I finished the shift with project 138-23, performing 2 more repetitions on NGC4214, after a pointing on 1144+402 (FWHM 16.73"x18.31"). Overall, very good stability, with <rms> < 40 mJy/beam.

06:00-17:50; Tuan-Anh, Quang, Ioannis, Angel (merged shift): starting with a calib_1scan scan (s35) on source MWC349 in total power. DAQ restarted in polarimetry mode (lost KIDs: 18, 2, 39). Getting pointing (redid after a hexapod error message in the ACU) and focusing first on K3-50A (nikaw-23 account). FWHM 17.5" x 19.4" after focus correction. Very stable atmosphere, tau_225Ghz < 0.1. Planned to observe DR12OH up to El. ~70-75 deg. and then Serp-SMM4. Switched to project 015-17, changed to 2200+420 (BL Lac) when its position was compliant with topo low without a round trip of the telescope. Getting a new pointing & focus correction (1.47 mm); FWHM 17.7" x 18.0" (2 mm). Got a first calib_1scan (scan 52). A total of 3 rounds on DR12OH (6 scans) with calibrations on BL Lac and 2 rounds on Serp-SMM4 (4 scans) with Mars for calib_1scan maps were obtained during the morning shift. Calibrations were launched between single or double rounds, depending on required pointing/focus corrections. Taking advantage of the superb conditions and the lower pressure of science targets in the afternoon, we extended the shift to get a new round set on DR12OH (6 scans). Switching to total power at 16:40; restarting DAQ (after checking twice anomalous behaviour of the board in slot 5, crate 1; no reset needed), lost KIDs: 26, 2, 43. Alternate fog around the telescope. Pointing on Uranus, the beam at 2 mm is a bit distorted (17.3" x 19.1"). No data are available from PIIC-monitors at 1 mm: imbFITS files of the preceding polarimetry session are being created at this time. No change of focus correction. New pointing check, beam geometry is improving: FWHM 17.4" x 18.3". Sofia takes over.

17:50-00:00; Sofia: I start with a pointing and focus on 0458-020 (FWHM 17.3" x 18.9") under nikaw-23. Focus correction 1.04mm. At 18:14 (local time) I switch to project 117-23, targeting orion b. Start with pointing on 0458-020, set pointing 5 9.8. I launch the first repetition of 117-23/observe-orionb-4. Launch a new pointing in nikaw-23 on 458-020. Set pointing 4.9 8, then I go to Uranus and get a new pointing. I set pointing correction 4.2 5.4, with FWHM 17.34" x 18.71". New focus on Uranus: not successful, wind speed ~ 12m/s. Switch to source 0316-413, very bright quasar. Trying to point and focus again. Not working. Switch to source 0954+658, so that the antenna is pointing north, opposite of the wind direction. Pointing on 0954+658. Tried pointing and focus, but it is still not working due to high wind speed (see plot below). Parking the telescope at 20h35 (local time), with current wind speed ~13 m/s. Wind forecasts are not promising for the next hours. <AB> Joaquín Santiago managed to record useful tracking data of 5 selected scans with Scope when wind speed was between ~10 m/s and 15 m/s. He stored the results in an ad-hoc ACU folder and sent related information to OHB, aimed at improving the behaviour of the tracking in these wind regimes. Thanks, Joaquín.


2024 Mar. 7th, Thursday

00:00-06:00; Frédéric: the wind peaks at ~20 m/s. The antenna is still parked.

06:00-14:00; Tuan-Anh, Quang: the wind peaks over 30 m/s, no observations.

14:00-18:00; Angel: average wind speed between ~15 and 30 m/s, dense fog; no observations.

18:00-00:00; Sofia: wind speed around 15 m/s. It started snowing. No observations.

2024 Mar. 8th, Friday

00:00-06:00; Frédéric: the wind speed still peaks around 10 m/s. No observations.

06:00-14:00; Tuan-Anh, Quang: telescope still parked because of ice and wind.

14:00-18:00; Angel: high wind speed and snow. No observations.

18:00-00:00; Sofia: high wind speed and snow. No observations.

2024 Mar. 9th, Saturday

00:00-06:00; Frédéric: high wind speed, strong gusts, snowing. No observations.

06:00-14:00; Tuan-Anh, Quang: telescope still parked because of ice and wind.

14:00-18:00; Angel: mean wind speed has dropped a bit, still snowing: increasing snow accumulation around. Deicing ~60%. No observations.

18:00-00:00; Sofia: No observations.

2024 Mar. 10th, Sunday

Snowfall at intervals, wind and fog all day. Deicing is still on. No observations.

Morning Veleta blues; one of the youngest members of the herd (courtesy of Mónica Rienda); snowstorm continues overnight:

attachment:30m_CR_20240310_11h12UT.jpg attachment:fox_1_30m_20240310.jpg attachment:30m_20240310_21h00UT.jpg

2024 Mar. 11th, Monday

00:00-06:00; Frédéric: light snow and wind. Clear skies at dawn. Telescope is frozen, deicing is still on. No observations.

06:00-14:00; Tuan-Anh, Quang: no wind, no clouds. Telescope defrosting is ongoing.

14:00-18:00; Angel: From the platform, Juan Luis cleaned the ice on the sub-reflector/rear support and the segments of quadrupod legs close to the secondary cell. Thanks, Juan Luis. Ice pieces remain on the outer ring of the dish and unreachable parts of the quadrupod. OHB tests on tracking/control loop are scheduled at 17h. Meanwhile, we hope that the remaining ice will loosen with the daytime heat/sunshine. Finally, the planned tests were not carried out because some pieces of ice prevented the telescope from moving in elevation due to the risk to the surface of the main dish. While the telescope was moved only in azimuth to facilitate the last phase of deicing, anomalous torque behaviour in motor group 2&4 caused the telescope to stop in safe mode from the ACU. After an initial inspection by technical and operations personnel, additional attempts to move the telescope in azimuth were affected by the same apparent failure. It was decided to keep the telescope stopped until further investigations are carried out, together with OHB technicians, during the next maintenance period.

18:00-00:00; Sofia: No observations.

2024 Mar. 12th, Tuesday

No observations.

DailyReportsNika2Pool050324 (last edited 2024-03-18 09:57:00 by NikaBolometer)